OpenStudy (anonymous):

solve each equation for 0≤x≤2π: sin x= √3 cos x explain why the trig equation is not an identity.

OpenStudy (anonymous):

first of all, the trig equation is not an identity because the value of 'x' can be found. (cosx)^2+(sinx)^2=1 is an identity because no matter what you do, the value of x cannot be found

OpenStudy (anonymous):

\[sinx=(\sqrt{3})cosx\]\[\frac{ sinx }{ cosx } = \sqrt{3}\]\[tanx=\sqrt{3}\]

OpenStudy (anonymous):

do you follow me?

OpenStudy (anonymous):

no Mousam :(

OpenStudy (anonymous):

oh... where did i lose you?

OpenStudy (anonymous):

the explanation or the working out?

OpenStudy (anonymous):

the working out

OpenStudy (anonymous):

oh ok.. do you get my explanation for why that is an equation, not an identity?

OpenStudy (anonymous):

ya i understand that. What i dont get is how is 3 square root between 0<2<pi?

OpenStudy (anonymous):

so your question is \[sinx=(\sqrt{3})cosx\]

OpenStudy (anonymous):

what does 0≤x≤2π mean?

OpenStudy (anonymous):

oh... you've misunderstood the question.... we have to find out values of X between 0<x<2pi

OpenStudy (anonymous):

woah... did you just start this topic in class?

OpenStudy (anonymous):

do you know what 2pi is?

OpenStudy (anonymous):

yaa

OpenStudy (anonymous):